OpenStudy (anonymous):

What is the best divisibility test for 11

OpenStudy (gorv):

abcdefg is number

OpenStudy (gorv):

now digit at even place are b d and f sum them up=b+d+f

OpenStudy (gorv):

sum of digit at odd place =a+c+e+g now difference of these to sum should be zero or divisible by 11

OpenStudy (gorv):

like 121

OpenStudy (gorv):

even place =2

OpenStudy (gorv):

odd place =1+1=2

OpenStudy (gorv):

difference=2-2=0

OpenStudy (gorv):

so it will be divisible by 11